OpenStudy (anonymous):

Help??? 3.A textbook company keeps track of its daily cost for printing whole textbooks to different schools, but can only print up to 500 books a day. The daily cost (C) to print x number of textbooks is given by C(x) = 3.50x + 1200. A.Using the function, determine one possible daily cost by evaluating the function for any input of your choice. B.For what value of x does C(x) = 1900? C.Interpret the statement C(20) = 1270. D.Interpret the statement C(w) = p. E.Explain whether C(40.5) = 1341.75 is correct and possible given the context of this problem.

OpenStudy (michele_laino):

please, for option A, you can evaluate the maximum daily cost, namely: C(500)=3.5*500+1200=... please complete

OpenStudy (michele_laino):

for option B. you have to solve this equation, please: \[3.5x+1200=1900\]

OpenStudy (michele_laino):

of course, you have to find the value of x

OpenStudy (anonymous):

How do I find the value of x? sorry I am just not very good with algebra

OpenStudy (michele_laino):

for option C. please note that: \[C(20)=3.5*20+1200=1270\]

OpenStudy (anonymous):

I am really confused sorry :/

OpenStudy (michele_laino):

please where is the problem?

OpenStudy (michele_laino):

how can I help you?

OpenStudy (michele_laino):

@MeganKat

OpenStudy (anonymous):

I find it all very confusing.